# Ultrafast optical excitation of magnons in 2D antiferromagnetic semiconductors via spin torque mediated by unbound electron-hole pairs and excitons: Signatures in magnonic charge pumping

### v4

@@ I Introduction
In turn, dynamics of $\mathbf{M}_{i}(t)$ modifies the quantum Hamiltonian [Eq. (11) in the SM] of electrons within the loop in Fig. 2. This establishes self-consistency between TDNEGF and LLG calculations, as the initially photoexcited spin current, $I^{S_{\alpha}}=\mathrm{Tr}[{\bm{\rho}}^{\rm neq}(t)\hat{I}^{S_{\alpha}}]$, will be modified by the dynamics of $\mathbf{M}_{i}(t)$; whereas their trajectories are, in turn, affected by updated spin current and STT $\mathbf{T}_{i}(t)$ exerted by it. In the LLG Eq. (3), $g_{0}$ is the gyromagnetic factor [68]; $\mu_{M}$ is the magnitude of LMMs [68]; $\mathbf{B}^{\rm eff}_{i}=-\frac{1}{\mu_{M}}\partial\mathcal{H}/\partial\mathbf{M}_{i}$ is the effective magnetic field obtained from classical Hamiltonian $\mathcal{H}$ [Eq. (10) in the SM] for LMMs [48, 50, 69]; Gilbert damping is chosen as $\alpha_{G}=0.01$, which is the typical value for Cr-based 2D magnets [50]; and we use shorthand notation $\partial_{t}\equiv\partial/\partial t$.
Figure 3: Time dependence of the Néel vector, defined [Eq. (8)] for two monolayers of AF semiconductor CAR [Fig. 1] in: (a) the absence of excitons ($U=0$); or (c) their presence induced by on-site Coulomb interaction [15, 19] $U=0.1\gamma$ [Eq. (14) in the SM]. Panels (b) and (d) plot the corresponding power spectrum of windowed FFT [90, 91], $|N^{z}(t_{c},\omega)|^{2}$, revealing frequencies and lifetimes of magnons excited by STT from the photocurrent of unbound electron-hole pairs in (a) or those pairs and excitons acting together in (b), respectively. Here $t_{c}$ denotes the central time of the Gaussian window [90, 91] used in FFT. Note that $\omega_{b}$ labels the frequency of the bright magnon, which corresponds to the same type of magnon observed in experiments of Refs. [3, 4, 5] on in bilayers of AF semiconductor CrSBr. Gray curves on the bottom of panels (a) and (c) depict the vector potential $A^{x}(t)$ of fsLP.
@@
for the time evolution of ${\bm{\rho}}^{\rm neq}(t)$, where $\mathbf{H}(t)$ is the matrix representation of the quantum Hamiltonian of electrons. Equation (4) is an exact quantum master equation for the reduced density matrix of the AF CAR in Fig. 2 viewed as an open finite-size quantum system attached to macroscopic Fermi liquid reservoirs via semi-infinite normal metal (NM) leads. The NM leads, not exploited in experiments [3, 4, 7, 8, 9, 5, 10, 11] thus far, are important technically within TDNEGF calculations in order to introduce continuous energy spectrum and dissipation effects  [92, 93], thereby guaranteeing that excited photocurrent and $\mathbf{M}_{i}(t)$ dynamics will eventually terminate at some time after fsLP cases. Otherwise, in a closed quantum system [89, 65] without a surrounding bath [73] one would find a forever oscillating photocurrent, which is unphysical. Furthermore, the NM leads allow us to analyze properties of charge and spin currents drained into them. Such currents could offer a novel experimental probe of excitons, magnons, and their interactions, as we confirm in Fig. 5. For this purpose, we use ${\bm{\Pi}}_{p}(t)$ matrices
@@
currents outflowing into $p=L,R$ NM leads. Since the applied bias voltage between the left ($L$) and right ($R$) NM leads is identically zero in the setup of Fig. 1, all computed currents $I_{p}(t)$ and $I_{p}^{S_{\alpha}}(t)$ are solely due to pumping [94, 95, 78, 80, 96, 52, 55] by (nonperiodic [94]) time-dependence of the CAR Hamiltonian. Note that we use the same units for charge and spin currents, defined as $I_{p}=I_{p}^{\uparrow}+I_{p}^{\downarrow}$ and $I_{p}^{S_{\alpha}}=I_{p}^{\uparrow}-I_{p}^{\downarrow}$, in terms of spin-resolved charge currents $I_{p}^{\sigma}$. In our convention, positive current in the NM lead $p$ means charge or spin current is flowing out of that NM lead.
@@
Let us recall that the problem of how STT excites uniform motion of all LMMs [i.e., of their macrospin $\mathbf{M}(t)=\sum_{i}\mathbf{M}_{i}(t)$] vs. their nonuniform motion like magnons was analyzed long ago [97] for conventional metallic ferromagnets, as well as recently for AF insulators [98, 81], with focus on threshold injected current value [43, 99] for magnons to occur. Our setup in Fig. 2 is different from those studies, as no current is injected from an external circuit. Instead, photocurrent is excited within the AF semiconductor CAR by fsLP, and can eventually exit into the NM leads. In addition to the photocurrent of conduction electrons, we also consider the photocurrent of excitons generated by Coulomb interaction, which binds conduction-band electrons with valence-band holes. To capture such binding, we employ a time-dependent mean-field theory (tMFT) [19, 15] of the inter-orbital Coulomb interaction term [Eq. (14) in the SM]. For this purpose, we exploit [19] the off-diagonal elements of $\mathbf{\bm{\rho}}^{\rm neq}(t)$, that we naturally construct within the TDNEGF part [76, 75] of the TDNEGF+LLG self-consistent loop [78, 79, 52, 80, 81]. In other words, for the study presented here, the previously developed [78, 79, 52, 80, 81, 100] TDNEGF+LLG framework is amended with modeling of excitonic effects into the TDNEGF+LLG+EX framework illustrated by Fig. 2.
Further details of our quantum Hamiltonian for the electronic subsystem, $\hat{H}$, and the classical one, $\mathcal{H}$, for the subsystem of LMMs, together with explanations of tMFT and electromagnetic (EM) radiation calculations via the Jefimenko equations [101, 102, 103, 104, 105, 94, 106], are provided in the SM.
II Results and Discussion
We first recall the definition [3, 4, 5, 9] of the Néel vector
@@ II Results and Discussion
between two monolayers of AF semiconductor, where in equilibrium $\mathbf{N}(t=0)\equiv(2,0,0)$. Out of equilibrium, as initiated by fsLP, the Néel vector starts evolving in time [Figs. 3(a) and 3(c)]. Somewhat surprisingly and also observed experimentally [14], this evolution starts without substantial delay regarding fsLP, even though LMMs are slower than electrons [73]. Because the magnon frequency spectrum encoded in $N^{z}(t)$ could be changing within different time frames, we apply windowed (or short-time) Fast Fourier transform (FFT) [90, 91, 107, 108, 109]. For example, magnons get excited around $t_{c}\simeq 15\ \mathrm{ps}$ in Fig. 3(b), and subsequently they decay [34, 37] because of Gilbert damping $\alpha_{G}$ in Eq. (3). Furthermore, because of explicitly introduced electrons via TDNEGF calculations, additional nonlocal damping [79, 110, 111, 112, 113, 114, 115] is introduced into Eq. (3) via the STT term $\mathbf{T}_{i}(t)$. Due to such decay, excited magnons vanish at around $t_{c}\simeq 18\ \mathrm{ps}$ in Fig. 3(b). For such signals—appearing also in many other scientific disciplines (such as electroencephalography [91] or speech analysis)—it is advantageous to perform windowed FFT over successive time intervals. For this purpose, we employ Gaussian as the window function
@@
which makes it possible to extract time-frequency content from signals whose oscillations are localized in a finite time frame. Here, $t_{c}$, the centroid of the Gaussian, serves as the abscissa of panels (b) and (d) within each of Figs. 35 analyzing $X(t)=N^{z}(t),I_{R}(t),E^{x}_{\mathrm{FF}}(t)$ as the signal, respectively. Here, $\sigma$ specifies the width of the Gaussian and $\lambda$ is a parameter controlling the resolution. For example, greater values of $\lambda\sigma$ yield better resolution in the time domain, while smaller values yield improved resolution in the frequency domain, where the frequency and time resolution satisfy a Heisenberg-like uncertainty relation [90, 91]. Windowed FFT of the Néel vector produces $N^{z}(\omega,t_{c})$, whose power spectrum in Figs. 3(b) and 3(d) reveals excitation of the so-called “bright magnon” at the frequency $\omega_{b}$. In the presence of excitons due to nonzero $U$, we find a longer lifetime of excited magnons in Fig. 3(d). Both Figs. 3(b) and  3(d) also show short (for $U=0$) vs. longer (for $U\neq 0$) living magnons, respectively, that are excited at frequencies below $\omega_{b}$. We confirm in Fig. 6(a) of the SM that such magnons are a direct consequence of $J_{K}\neq 0$. Thus, our TDNEGF+LLG+EX framework explains microscopically the same “bright magnon” observed in experiments on CrSBr [3, 4, 5, 9] using fsLP whose central frequency is above-gap as the consequence of time-dependent STT [Eq. (1)] mediated by photoexcited unbound electron-hole pairs and excitons.
@@
Figure 5: Time-dependence of the $E_{\mathrm{FF}}^{x}(t)$ component of the electric field of EM radiation emitted into the FF region by magnon (from Fig. 3) generated bond charge currents [Eq. (17) in the SM] within the CAR of Fig. 1 in: (a) the absence of excitons ($U=0$); or (c) their presence induced by on-site Coulomb interaction $U=0.1\gamma$. Panels (b) and (d) plot the corresponding power spectrum of windowed FFT [90, 91], $|E_{\mathrm{FF}}^{x}(t_{c},\omega)|^{2}$. Gray curves on the bottom of panels (a) and (c) depict the vector potential $A^{x}(t)$ of fsLP.
@@
The excited magnons will introduce the second nonequilibrium drive into the subsystem of electrons. Such a time-dependent drive can lead to pumping [53, 52, 54, 55] of electronic spin and charge currents. They can be differentiated from currents pumped [82] by fsLP (as the first nonequilibrium drive) by their spectral content in the frequency domain. We compute charge current $I_{R}(t)$ [Figs. 4(a) and 4(c)] pumped into the right NM lead, and perform windowed FFT on it to obtain $|I_{R}(t_{c},\omega)|^{2}$ plotted [Figs. 4(b) and 4(d)] in the same frequency range where magnons are found in Fig. 3. The power spectrum $|I_{R}(t_{c},\omega)|^{2}$ exhibits a peak at $\omega_{b}$, thus, by attaching an additional external circuit to recent experiments and by analyzing pumped current into that circuit, one could confirm optical excitations of magnons. Interestingly, pumped current could also sense the presence of excitons—in $U\neq 0$ case long-lived high harmonic generation (HHG) emerges in Fig. 4(d) particularly at the frequency $4\omega_{b}$.
@@
Finally, we examine EM radiation emitted by time-dependent local (or bond [116]) charge currents $I_{ia\rightarrow jb}(t)$ [Eq. (17) in the SM] within AF semiconductor CAR in Fig. 1. Note that EM radiation emitted by ultrafast-light-driven magnetic materials and their heterostructures is routinely used [117, 118, 119, 120] in spintronic experiments as a probe of coupled spin-charge dynamics [105, 121, 94, 106] of such far-from-equilibrium systems. We compute the $x$-component $E^{x}_{\mathrm{FF}}(\mathbf{r},t)$ of the electric field of EM radiation in the far-field (FF) region via the Jefimenko [101, 103] formula using $I_{ia\rightarrow jb}(t)$ as the source [104, 73]. The FF region is defined by radiation decaying as $\sim 1/r$, so we use $\mathbf{r}_{0}=(5a,0,1000a)$ as the observation point in the Jefimenko formula [Eq. (16) in the SM]. The windowed FFT of $E^{x}_{\mathrm{FF}}(\mathbf{r}_{0},t)$ in Figs. 5(a) and 5(c) yields $E^{x}_{\mathrm{FF}}(\mathbf{r}_{0},t_{c},\omega)$ whose power spectrum is plotted in Figs. 5(b) and 5(d), respectively. Similarly to $|I_{R}(t_{c},\omega)|^{2}$, the power spectrum $|E^{x}_{\rm FF}(t_{c},\omega)|^{2}$ contains an imprint of excited “bright magnon” at $\omega_{b}$ frequency, as well as HHG at $4\omega_{b}$ [Fig. 5(d)] signaling the presence of excitons and their role in STT exciting magnons in Fig. 3. Note that the frequency content of EM radiation by time-dependent current generated by excited magnons will be in the range of: THz in Fig. 5 due to parameters selected in our quantum and classical Hamiltonians for convenience (i.e., to ensure we can complete time evolution within $\sim 10$ ps); sub-THz for realistic CrI3 in Ref. [14]; and GHz for realistic CrSBr in Ref. [5].
III Conclusions and Outlook

### v5

@@ I Introduction
In turn, dynamics of $\mathbf{M}_{i}(t)$ modifies the quantum Hamiltonian [Eq. (6)] of electrons within the loop in Fig. 2. This establishes self-consistency between TDNEGF and LLG calculations, as the initially photoexcited spin current, $I^{S_{\alpha}}=\mathrm{Tr}[{\bm{\rho}}^{\rm neq}(t)\hat{I}^{S_{\alpha}}]$, will be modified by the dynamics of $\mathbf{M}_{i}(t)$; whereas their trajectories are, in turn, affected by updated spin current and STT $\mathbf{T}_{i}(t)$ exerted by it. In the LLG Eq. (4), $g_{0}$ is the gyromagnetic factor [59]; $\mu_{M}$ is the magnitude of LMMs [59]; $\mathbf{B}^{\rm eff}_{i}=-\frac{1}{\mu_{M}}\partial\mathcal{H}/\partial\mathbf{M}_{i}$ is the effective magnetic field obtained from classical Hamiltonian $\mathcal{H}$ [Eq. (5)] for LMMs [48, 50, 60]; Gilbert damping is chosen as $\alpha_{G}=0.01$, which is the typical value for Cr-based 2D magnets [50]; and we use shorthand notation $\partial_{t}\equiv\partial/\partial t$.
The paper is organized as follows. The quantum Hamiltonian for the electronic subsystem, $\hat{H}$, and the classical one, $\mathcal{H}$, for the subsystem of LMMs are introduced in Sec. II.1. The TDNEGF+LLG+EX methodology, using these Hamiltonians as an input, is covered in Sec. II.2, as well as visually in Fig. 2. Section II.3 explaines the algorithm for the computation of electromagnetic (EM) radiation via the Jefimenko equations [92, 93, 94, 95, 96, 97, 98], which uses bond currents as an input obtained from time-dependent quantum transport simulations via the TDNEGF+LLG+EX framework. We conclude and provide an outlook in Sec. IV. In addition, we provide Fig. 6 in the Appendix—with parameters from Fig. 3 being artificially switched off or substantially increased, which then helps to further elucidate its results—as well as a movie as the Supplemental Material (SM)  [^1] See Supplemental Material at https://wiki.physics.udel.edu/qttg/Publications for a movie animating the dynamics of LMMs and STT exerted on them, as well as charge currents pumped by dynamical LMMs (i.e., by magnons as a collective dynamics of many LMMs). animating Figs. 3 and  4.
II Models and Methods
II.1 Quantum and classical Hamiltonians
Our TDNEGF+LLG+EX calculations are fully microscopic, requiring only two Hamiltonians as an input. These two Hamiltonians describe bare degrees of freedom—quantum Hamiltonian, $\hat{H}(t)$, for electrons; and the classical one, $\mathcal{H}(t)$, for LMMs. In this study, we focus on essential features of bilayer 2D AF semiconductors [12, 11, 63] by constructing a model quantum Hamiltonian [Eq. (6)] that captures two monolayers, a semiconducting bandgap for each of them, ferromagnetic (FM) intralayer and AF interlayer ordering of LMMs, and possible Coulomb interaction effects [15, 19] binding conduction-band electrons and valence-band holes into excitons. Accordingly, our simulation setup in Fig. 1 consists of two one-dimensional (1D) tight-binding (TB) chains [46, 47, 51] with two electronic orbitals per site $i$, as well as one LMM per site $i$ described by the unit vector $\mathbf{M}_{i}(t)$. Our usage of 1D chains is inspired by CrSBr [46, 47, 48, 49, 50, 51] as a highly anisotropic 2D magnetic material [1, 2] formed [46, 47, 51] by 1D atomic chains with interlayer AF ordering (A-type) and intralayer FM ordering. The Curie temperature of FM ordering within a single monolayer is $\simeq 150$ K, while AF ordering of LMMs between adjacent layers has a Néel temperature of $\simeq 130$ K [49]. Note that angle-resolved photoemission spectroscopy reveals the complex electronic band structure of CrSBr, whose accurate theoretical description requires first-principles methods, such as DFT+U [51] or self-consistent GW methodology [49, 10, 23].
The classical Hamiltonian, entering into the LLG Eq. (4), is given by
Such an effective classical Heisenberg Hamiltonian of LMMs can be extracted from first-principles calculations [100, 74, 101] or by fitting magnon spectra from neutron scattering data [48]. Here $J^{\rm AF}=0.0195\ \mathrm{eV}$ and $J^{\rm FM}=0.15\ \mathrm{eV}$ are AF and FM exchange couplings, respectively; $K_{x}=0.021\ \mathrm{eV}$ and $K_{z}=0.057\ \mathrm{eV}$ specify the magnetic anisotropy along the $x$- and $z$-axis, respectively; $J_{K}=0.01\ \mathrm{eV}$ is Kondo exchange interaction between spins of flowing electrons and LMMs [88, 89]; $i\in 1$ ($i\in 2$) means that the site $i$ belongs to the first (second) layer in Fig. 1; and $\langle\ldots\rangle$ denotes NN sites. Note that $J^{\rm AF},\ J^{\rm FM},\ K_{x},\ K_{z}$ parameters were also employed to interpret experiments [3, 4, 5], but we rescale their values by a multiplicative factor $\times 10^{3}$ to make the total TDNEGF+LLG+EX simulation time of $\sim 10$ ps duration. Nevertheless, even with this adjustment, the characteristic energy scale of the AF background remains below electron kinetic energy, $\gamma_{0}/J^{\rm{AF}}\simeq 10^{2}$. Such rescaling shifts the spectral content of pumped charge current and its radiation into the THz range in Figs. 4 and  5, while for, e.g., CrSBr the relevant range would be at GHz frequencies.
Here $\hat{H}_{\mathrm{intra}}$ describes 1D TB chains within layer 1 or 2 in Fig. 1
$\hat{H}_{\mathrm{intra}}=\Delta/2\sum_{i\in 1,2}(\hat{c}_{i}^{c\dagger}\hat{c}^{c}_{i}-\hat{c}_{i}^{v\dagger}\hat{c}^{v}_{i})$ (7)
where indices $c$ and $v$ stand for orbitals that give rise to the conduction and valence bands of each chain; $\hat{c}_{i}^{a\dagger}=(\hat{c}_{i\uparrow}^{a\dagger}\ \ \hat{c}_{i\downarrow}^{a\dagger})$ is the row vector containing operators $\hat{c}_{i\sigma}^{a\dagger}$ creating electrons with spin $\sigma=\uparrow,\downarrow$ in orbital $a=c,v$ hosted by site $i$; $\Delta$ is the onsite potential opening a bandgap between the two bands $\Delta=3\ \rm{eV}$; and $\gamma_{0}=1\ \mathrm{eV}$ is the hopping parameter between the NN sites. The spin density operator in Eqs. (3) and  (6) is given by $\mathbf{\hat{s}}_{ia}=\hat{c}_{i}^{a\dagger}\hat{\bm{\sigma}}\hat{c}^{a}_{i}$ for $a=c,v$ and ${\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ being the vector of the Pauli matrices. The fsLP is introduced in Eq. (7) via the Peierls phase [102, 103], $\chi_{ij}(t)=z_{\max}\exp[-\left(t-t_{p}\right)^{2}/(2\sigma_{\text{light }}^{2})]\sin(\omega_{0}t)$, where the electric field of the pulse is $\mathbf{E}(t)=-\partial_{t}\mathbf{A}(t)$, $z_{\max}=eaA_{\rm max}/\hbar=0.1$ is a dimensionless parameter quantifying the intensity of the pulse using $A_{\rm max}$ as the amplitude of the vector potential, $\omega_{0}$ is the central frequency of fsLP, and $a$ is the lattice spacing. Besides the Peierls phase, fsLP is additionally [15] introduced via hopping $\gamma_{\mathrm{P}}(t)=\mathbf{d}\cdot\mathbf{E}(t)$ in Eq. (7). This term accounts for interband transitions driven by the dipole interaction with the electric field [15], where $\mathbf{d}$ is the expectation value of the dipole operator, $\mathbf{d}=e\langle i,c|\mathbf{\hat{r}}|i,v\rangle$. The term $\hat{H}_{\mathrm{inter}}$ is given by
and it describes hopping with parameters $\gamma_{c}=0.5\gamma_{0}$ or $\gamma_{v}=0.5\gamma_{0}$ between $c$ or $v$ orbitals, respectively, located at NN sites of two different chains. Finally, the inter-orbital and intralayer Coulomb interaction [15, 19]
describes how two electrons on two different orbitals at the same site $i$ interact with each other. The same Hubbard term was employed in prior studies [15, 19] to describe exciton formation on the TB lattice. We decouple it via tMFT [19] (otherwise, its beyond-tMFT treatment requires computationally much more expensive evaluation of Feynman diagrams on the Keldysh contour [15]) as follows:
Here the first two terms correspond to Hartree and the latter two to Fock approximation, where the order parameter of the excitonic condensate [36] is given by $\phi_{i}(t)=\langle\hat{c}_{i}^{v\dagger}(t)\hat{c}^{c}_{i}(t)\rangle$. This procedure requires us to self-consistently compute $\phi_{i}(t)$, which we obtain from the off-diagonal elements of ${\bm{\rho}}^{\rm neq}(t)$ as a part of TDNEGF calculations within the self-consistent loop illustrated in Fig. 2. We note that in this approach we do not rely on the effective excitonic Hamiltonians written in terms of operators of composite bosons [73], which can be cumbersome to derive [104]. Instead, we use “bare” degrees of freedom, so that both unbound electron-hole pairs and their binding into excitons by the Coulomb interaction are considered through the electronic density matrix, as naturally included [19, 15, 16, 17] when using the nonequilibrium density matrix and/or TDNEGF formalism.
The AF semiconductor CAR in Fig. 1 is attached to $L$ and $R$ normal metal (NM) leads modeled as semi-infinite ideal 1D TB chains with one orbital per site. The chemical potential of the macroscopic reservoirs into which NM leads terminate is identical (i.e., no DC bias voltage is applied) and chosen as $\mu_{L}=\mu_{R}=E_{F}=0$.
II.2 TDNEGF+LLG+EX methodology
@@
for the time evolution of ${\bm{\rho}}^{\rm neq}(t)$, where $\mathbf{H}(t)$ is the matrix representation of the quantum Hamiltonian of electrons. Equation (11) is an exact quantum master equation for the reduced density matrix of the AF CAR in Fig. 2 viewed as an open finite-size quantum system attached to macroscopic Fermi liquid reservoirs via semi-infinite NM leads. The NM leads, not exploited in experiments [3, 4, 7, 8, 9, 5, 10, 11] thus far, are important technically within TDNEGF calculations to introduce continuous energy spectrum and dissipation effects  [105, 106], thereby guaranteeing that excited photocurrent and $\mathbf{M}_{i}(t)$ dynamics will eventually terminate [80] after fsLP cases. Otherwise, in a closed quantum system [91, 73] without a surrounding bath [80] one would find a forever oscillating photocurrent (as it occurs, e.g., in first-principles time-dependent density functional theory calculations applied to closed quantum systems [96, 107, 108]), which is unphysical. Furthermore, the NM leads allow us to analyze properties of charge and spin currents drained into them. Such currents could offer a novel experimental probe of excitons, magnons, and their interactions, as we confirm in Fig. 4. For this purpose, we use ${\bm{\Pi}}_{p}(t)$ matrices
@@
currents outflowing into $p=L,R$ NM leads. Since the applied bias voltage between the left ($L$) and right ($R$) NM leads is identically zero in the setup of Fig. 1, all computed currents $I_{p}(t)$ and $I_{p}^{S_{\alpha}}(t)$ are solely due to pumping by nonperiodic [97, 109, 61, 82] (instead of usually considered periodic [97, 110, 52, 55]) time dependence of the CAR Hamiltonian. Note that we use the same units for charge and spin currents, defined as $I_{p}=I_{p}^{\uparrow}+I_{p}^{\downarrow}$ and $I_{p}^{S_{\alpha}}=I_{p}^{\uparrow}-I_{p}^{\downarrow}$, in terms of spin-resolved charge currents $I_{p}^{\sigma}$. In our convention, positive current in the NM lead $p$ means charge or spin current is flowing out of that NM lead.
@@
Let us recall that the problem of how STT excites uniform motion of all LMMs [i.e., of their macrospin $\mathbf{M}(t)=\sum_{i}\mathbf{M}_{i}(t)$] vs. their nonuniform motion like magnons was analyzed long ago [111] for conventional metallic ferromagnets, as well as recently for AF insulators [112, 83], with a focus on threshold injected current value [43, 113] for magnons to occur. Our setup in Fig. 2 is different from those studies, as no current is injected from an external circuit. Instead, photocurrent is excited within the AF semiconductor CAR by fsLP. In addition to the photocurrent of unbound electron-hole pairs, we also consider possibility of photocurrent of excitons due to Coulomb interaction [Eq. (9)] binding conduction-band electrons with valence-band holes. To capture such binding, we employ a time-dependent mean-field theory (tMFT) [19, 15] of the inter-orbital Coulomb interaction term [Eq. (9)]. For this purpose, we exploit [19] the off-diagonal elements of $\mathbf{\bm{\rho}}^{\rm neq}(t)$, that we naturally construct within the TDNEGF part [57, 56] of the TDNEGF+LLG self-consistent loop [61, 62, 52, 82, 83]. In other words, for the study conducted in Sec. III, the previously developed [61, 62, 52, 82, 83, 114] TDNEGF+LLG is generalized into the TDNEGF+LLG+EX framework by adding modeling of excitonic effects at tMFT level [19, 15], as illustrated by magenta box in Fig. 2.
II.3 Electromagnetic radiation by bond charge currents from the Jefimenko equations
Note that Jefimenko Eq. (15) can be viewed [93] as proper (i.e., time-retarded) time-dependent generalizations of the Coulomb law. Here, $t_{r}\equiv t-|\mathbf{r}-\mathbf{l}|/c$ emphasizes retardation in the response time due to relativistic causality [92, 94]. Additionally, we adapt [95, 80] Eq. (15) to utilize time-dependent bond [115] charge currents as the source of EM radiation, $I_{ia\rightarrow jb}(t)$ [Eq. (16)]. They are the counterpart (on the TB lattice) of local current density in continuous space. The bond currents $I_{ia\rightarrow jb}$ are assumed to be spatially homogeneous along the path $P_{ia\rightarrow jb}$ from orbital $a$ at site $i$ to orbital $b$ at site $j$ [115, 61, 95], which is composed of a set of points $l\in P_{ia\rightarrow jb}$. Here $N_{B}$ is the number of bonds $ia\rightarrow jb$, and since we use $N=10$ black and gray sites [Fig. 1] in each layer of CrSBr, $N_{B}=36$ in our calculations. We obtain bond charge currents as
III Results and Discussion
Figure 3: Time dependence of the Néel vector, defined [Eq. (17)] for two monolayers of 2D AF semiconductor in Fig. 1 in: (a) the absence of excitons ($U=0$); or (c) their presence induced by on-site Coulomb interaction [15, 19] $U=0.1\gamma$ [Eq. (9)]. Panels (b) and (d) plot the corresponding power spectrum of windowed FFT [116, 117], $|N^{z}(t_{c},\omega)|^{2}$, revealing frequencies and lifetimes of magnons excited by STT from the photocurrent of unbound electron-hole pairs in (a) or those pairs and excitons acting together in (b), respectively. Here $t_{c}$ denotes the central time of the Gaussian window [116, 117] used in FFT [Eq. (18)]. Note that $\omega_{b}$ labels the frequency of the “bright magnon,” which corresponds to the same type of magnon observed in experiments of Refs. [3, 4, 5] on a bilayer of 2D AF semiconductor CrSBr. Gray curves on the bottom of panels (a) and (c) depict the vector potential $A^{x}(t)$ of fsLP.
We first recall the definition of the Néel vector analyzed in recent experiments [3, 4, 5, 9]
@@ III Results and Discussion
between two monolayers of AF semiconductor. Thus, in equilibrium and without any canting external magnetic field applied, we find $\mathbf{N}(t=0)\equiv(2,0,0)$. Out of equilibrium, as initiated by fsLP, the Néel vector starts evolving in time [Figs. 3(a) and 3(c)]. Somewhat surprisingly and also observed experimentally [14], this evolution starts without substantial delay regarding fsLP, even though LMMs are slower than electrons [80]. Because the magnon frequency spectrum encoded in $N^{z}(t)$ could be changing within different time frames, we apply windowed (or short-time) Fast Fourier transform (FFT) [116, 117, 118, 119, 120]. For example, magnons get excited around $t_{c}\simeq 15\ \mathrm{ps}$ in Fig. 3(b), and subsequently they decay [34, 37] because of Gilbert damping $\alpha_{G}$ in Eq. (4). Furthermore, because of explicitly introduced electrons via TDNEGF calculations, additional nonlocal damping [62, 121, 122, 123, 124, 125, 126] is introduced into Eq. (4) via the STT term $\mathbf{T}_{i}(t)$. Due to such decay, excited magnons vanish at around $t_{c}\simeq 18\ \mathrm{ps}$ in Fig. 3(b).
For such signals—appearing also in many other scientific disciplines (such as electroencephalography [117] or speech analysis)—it is advantageous to perform windowed FFT over successive time intervals. For this purpose, we employ Gaussian as the window function
@@
which makes it possible to extract time-frequency content from signals whose oscillations are localized in a finite time frame. Here, $t_{c}$, the centroid of the Gaussian, serves as the abscissa of panels (b) and (d) within each of Figs. 35 analyzing $X(t)=N^{z}(t),I_{R}(t),E^{x}_{\mathrm{FF}}(t)$ as the signal, respectively. Here, $\sigma$ specifies the width of the Gaussian and $\lambda$ is a parameter controlling the resolution. For example, greater values of $\lambda\sigma$ yield better resolution in the time domain, while smaller values yield improved resolution in the frequency domain, where the frequency and time resolution satisfy a Heisenberg-like uncertainty relation [116, 117]. Windowed FFT of the Néel vector produces $N^{z}(\omega,t_{c})$, whose power spectrum in Figs. 3(b) and 3(d) reveals excitation of the so-called “bright magnon”  [^2] The jargon “bright magnon” was introduced in experiments on CrSBr [3, 4, 5, 9]. For a clear explanation of the distinction between optically bright and dark magnon modes, see Ref. [4]. at the frequency $\omega_{b}$. In the presence of nonzero $U$ and thereby generated excitons, we find a longer lifetime of excited magnons in Fig. 3(d). Interestingly, both Figs. 3(b) and  3(d) show short (for $U=0$) vs. longer (for $U\neq 0$) living magnons, respectively, that are excited at frequencies below $\omega_{b}$. We, therefore, provide additional analysis in Appendix A confirming [Fig. 6(a)] that such magnons are a direct consequence of $J_{K}\neq 0$. Thus, our TDNEGF+LLG+EX framework explains microscopically the same “bright magnon” observed in experiments on CrSBr [3, 4, 5, 9], using fsLP whose central frequency is above-gap, as the consequence of time-dependent STT [Eq. (1)] mediated by photoexcited unbound electron-hole pairs or both them and excitons.
@@
Figure 5: Time-dependence of the $E_{\mathrm{FF}}^{x}(t)$ component of the electric field of EM radiation emitted into the FF region by excited “bright magnon” (from Fig. 3) pumping [53, 52, 54, 55] bond charge currents [Eq. (16)] within the CAR of Fig. 1 in: (a) the absence of excitons ($U=0$); or (c) their presence induced by on-site Coulomb interaction $U=0.1\gamma$. Panels (b) and (d) plot the corresponding power spectrum of windowed FFT [116, 117], $|E_{\mathrm{FF}}^{x}(t_{c},\omega)|^{2}$. Gray curves on the bottom of panels (a) and (c) depict the vector potential $A^{x}(t)$ of fsLP.
@@
Furthermore, the excited magnons will introduce the second nonequilibrium drive into the subsystem of electrons. Such a time-dependent drive can lead to pumping [53, 52, 54, 55] of electronic spin and charge currents. They can be differentiated from currents pumped [84] by fsLP (as the first nonequilibrium drive) by their spectral content, as fsLP and magnons have vastly different frequencies. We compute charge current $I_{R}(t)$ [Figs. 4(a) and 4(c)] pumped into the right NM lead, and perform windowed FFT on it to obtain $|I_{R}(t_{c},\omega)|^{2}$ plotted [Figs. 4(b) and 4(d)] in the same frequency range where magnons are found in Fig. 3. The power spectrum $|I_{R}(t_{c},\omega)|^{2}$ exhibits a peak at $\omega_{b}$, thus, by attaching an additional external circuit to recent experiments and by analyzing pumped current into that circuit, one could confirm optical excitations of magnons. Interestingly, pumped current could also sense the presence of excitons—in $U\neq 0$ case long-lived high harmonic generation (HHG) emerges in Fig. 4(d), particularly at frequency $4\omega_{b}$.
@@
Finally, we examine EM radiation emitted by time-dependent local (or bond [115]) charge currents $I_{ia\rightarrow jb}(t)$ [Eq. (16)] within AF semiconductor CAR in Fig. 1. Note that EM radiation emitted by ultrafast-light-driven magnetic materials and their heterostructures is routinely used [128, 129, 130, 131] in spintronic experiments as a probe of coupled spin-charge dynamics [96, 107, 97, 98] of such far-from-equilibrium systems. We compute the $x$-component $E^{x}_{\mathrm{FF}}(\mathbf{r},t)$ of the electric field of EM radiation in the far-field (FF) region via the Jefimenko [92, 94] formula using $I_{ia\rightarrow jb}(t)$ as the source [95, 80]. The FF region is defined by radiation decaying as $\sim 1/r$, so we use $\mathbf{r}_{0}=(5a,0,1000a)$ as the observation point in the Jefimenko formula [Eq. (15)]. The windowed FFT of $E^{x}_{\mathrm{FF}}(\mathbf{r}_{0},t)$ in Figs. 5(a) and 5(c) yields $E^{x}_{\mathrm{FF}}(\mathbf{r}_{0},t_{c},\omega)$ whose power spectrum is plotted in Figs. 5(b) and 5(d), respectively. Similarly to $|I_{R}(t_{c},\omega)|^{2}$, the power spectrum $|E^{x}_{\rm FF}(t_{c},\omega)|^{2}$ contains an imprint of excited “bright magnon” at $\omega_{b}$ frequency, as well as HHG at $4\omega_{b}$ [Fig. 5(d)] signaling the presence of excitons and their role in STT exciting magnons in Fig. 3. Note that the frequency content of EM radiation by time-dependent current generated by excited magnons will be in the range of: THz in Fig. 5 due to parameters selected in our quantum and classical Hamiltonians for convenience (i.e., to ensure we can complete time evolution within $\sim 10$ ps); sub-THz for realistic CrI3 in Ref. [14]; and GHz for realistic CrSBr in Ref. [5].
IV Conclusions and Outlook
IV.1 Conclusions